# When is Truthfully Allocating Chores no Harder than Goods?

### v1

@@ Abstract
For indivisible chores with two agents, we observe that a simple bundle-swapping operation transforms several properties for goods including truthfulness to the corresponding properties for chores, which enables us to characterize truthful mechanisms and derive the tight guarantees of various fairness notions achieved by truthful mechanisms. Moreover, for divisible chores, by generalizing the above transformation to an arbitrary number of agents, we characterize truthful mechanisms with two agents, show that every truthful mechanism with two agents admits an efficiency ratio of 00, and derive a large family of strictly truthful, envy-free (EF), and proportional mechanisms for an arbitrary number of agents. Finally, for indivisible chores with an arbitrary number of agents having bi-valued cost functions, we give an ex-ante truthful, ex-ante Pareto optimal, ex-ante EF, and ex-post envy-free up to one item mechanism, improving the best guarantees for bi-valued instances by prior works.
@@ 1 Introduction
When items are indivisible, [ABCM17] characterize the entire family of truthful mechanisms for two agents with additive valuations, which implies that truthfulness is incompatible with any non-trivial fairness notion. Nevertheless, this impossibility result can be bypassed by studying binary valuations [HPPS20, BEF21, BV22], allowing monetary transfer [GIK+24], or resorting to randomized mechanisms [BT24].
For (homogeneous) divisible items, [CGG13a] give a truthful mechanism that achieves a $1/e$-approximation of Nash welfare with some items unallocated. More recently, [FWVP23] characterize the set of truthful mechanisms for two agents (subject to some mild technical conditions) and provide a large family of strictly truthful, PROP, and EF mechanisms.
@@
In stark contrast to the fruitful outcomes for truthful good allocation, our understanding for truthful chore allocation still remains considerably more limited, and non-trivial results are only known for indivisible items. In particular, [ALW24] study a randomized mechanism that is truthful in expectation and achieves an expected approximation ratio of $O(\sqrt{\log n})$ for MMS. Moreover, [SC24] give a fair and efficient randomized mechanism for restricted additive valuations [^2] Restricted additive valuations constitute a strict superset of binary valuations and a strict subset of additive valuations., whose truthfulness also holds in expectation.
@@ 1.1 Our Contributions
For indivisible items with two agents, we show that a simple operation is able to transform numerous properties of a mechanism for goods to the counterparts of these properties for chores: Given a mechanism with desirable guarantees for goods, we invoke it with agents’ cost functions as input to obtain an allocation, and we output the resulting allocation by swapping agents’ bundles (Theorem 3.1). This, together with the characterization by [ABCM17] for goods, allows us to characterize the family of truthful mechanisms for chores with two agents (Theorem 3.2) and derive the tight guarantees achieved by truthful mechanisms with chores for various fairness notions (Corollary 3.3 and Corollary 3.4).
@@
Moving to divisible items, we notice that the above reduction from chore allocation to good allocation with two agents can be generalized to an arbitrary number of agents. To illustrate, suppose that we are given a mechanism with prominent properties for goods, and we invoke it with agents’ cost functions as input to obtain an allocation. Intuitively, in this allocation, each agent’s bundle consists of items that are highly undesired by him, and hence, we uniformly distribute each agent’s bundle among all other agents. We show that such an operation transforms fairness and truthfulness for goods to the corresponding properties for chores (Theorem 4.1), which, combined with the results by [FWVP23], enables us to characterize the family of truthful mechanisms for chores with two agents subject to some mild technical conditions (Theorem 4.6) and show that the efficiency ratio, the multiplicative loss in social welfare, of every truthful mechanism for chores with two agents is 00 (Theorem 4.7). Moreover, the family of strictly truthful, PROP, and EF mechanisms for goods with an arbitrary number of agents given by [FWVP23] can be easily adapted to mechanisms for chores with the same guarantees (see the remarks after Theorem 4.1).
@@
Finally, we study indivisible items for an arbitrary number of agents with bi-valued cost functions, i.e., the cost of each agent for each item is either $p$ or $q$ for some $p>q>0$, and the bi-valued setting has received extensive attention in recent years [ABF+21a, GM23, EPS22, GMQ22]. Recall that a randomized mechanism satisfies a property ex-ante if it holds in expectation, and satisfies a property ex-post if it holds for every realized integral allocation. We give a randomized mechanism that satisfies ex-ante truthfulness, ex-ante Pareto optimality (PO), ex-ante EF, and ex-post EF1 for chores (Theorem 5.1). This strengthens the guarantees of mechanisms by prior works [EPS22, GMQ22] for the same setting, which only satisfy ex-post fractional PO (fPO) and ex-post EF1 for chores. [^3] Note that ex-ante PO implies ex-post fPO. To achieve this, we leverage the mechanism for divisible goods with bi-valued agents recently introduced by [BT24] and the implementation scheme of [Azi20]. In particular, we largely exploit the balance property of the allocations returned by both of the aforementioned components, and considerable efforts are needed to prove the promised guarantees of our mechanism.
@@ 1.2 Other Related Work
During the past decades, there have been plentiful results in the area of fair allocation. In this subsection, we only focus on the most related works in the strategic setting and refer to the recent surveys [ABFV22, ALMW22, AAB+23] for a more comprehensive overview of the fair allocation literature.
@@
Besides the cases where resources are modeled as indivisible or divisible homogeneous items, another popular model assumes resources to be a heterogeneous divisible item, which is often referred to as cake cutting. The first truthful and EF cake-cutting mechanism for piecewise-uniform valuations is given by [CLPP13], which is later shown to be equivalent to the Maximum Nash Welfare mechanism and satisfy a stronger notion of group strategyproofness [AY14]. For the more general piecewise-constant valuations, [BST23] show the non-existence of truthful and PROP cake-cutting mechanisms even for two agents.
@@ 2 Preliminaries
Let $N=[n]$ denote the set of agents and $O=\{o_{1},\ldots,o_{m}\}$ denote the set of items. We will assume that either all items are goods or all of them are chores, and use utilities and costs to respectively describe the values of items. In particular, goods bring non-negative utilities, and agents are happier with a higher utility; conversely, chores bring non-negative costs, and agents are happier with a lower cost. We define the settings of indivisible items and divisible items separately.
@@ 2.1 Indivisible Items
2.2 Divisible Items
For divisible items, a (fractional) allocation $\mathbf{x}=(x_{1},\ldots,x_{n})\in[0,1]^{n\times m}$ satisfies $\sum_{i\in N}x_{i}(o)=1$ for every $o\in O$, where $x_{i}$ denotes the bundle received by agent $i$ and $x_{i}(o)$ denotes the fraction of item $o$ allocated to agent $i$. When items are goods, each agent $i$ is associated with a non-negative utility function $v_{i}$ where $v_{i}(x_{i})$ represents the utility of agent $i$ for every bundle $x_{i}$. For every $S\subseteq O$, we slightly abuse notations and use $v_{i}(S)$ to denote $v_{i}(\mathbf{1}_{S})$, where $\mathbf{1}_{S}\in[0,1]^{m}$ is the bundle such that $\mathbf{1}_{S}(o)=1$ for every $o\in S$ and $\mathbf{1}_{S}(o)=0$ for every $o\in O\setminus S$. For each $o\in O$, we use $v_{i}(o)$ to denote $v_{i}(\mathbf{1}_{\{o\}})$. We assume utility functions to be additive, i.e., $v_{i}(x_{i})=\sum_{o\in O}x_{i}(o)\cdot v_{i}(o)$. We normalize agents’ utility functions so that $v_{i}(O)=1$ for every $i\in N$. We define the cost functions $c_{i}(\cdot)$ in the same manner for chores. Given an allocation $\mathbf{x}$, the utility (resp. cost) of each agent $i$ is defined as $v_{i}(x_{i})$ (resp. $c_{i}(x_{i})$).
@@ 2.3 Mechanisms
A mechanism $\mathcal{M}$ is truthful for goods (resp. chores) if for every utility profile $\mathbf{v}=(v_{1},\ldots,v_{n})$ (resp. cost profile $\boldsymbol{c}=(c_{1},\ldots,c_{n})$), for all $i\in N$ and utility function $v_{i}^{\prime}$ (resp. cost function $c_{i}^{\prime}$), $v_{i}(\mathcal{M}_{i}(\mathbf{v}))\geq v_{i}(\mathcal{M}_{i}(v_{i}^{\prime},v_{-i}))$ (resp. $c_{i}(\mathcal{M}_{i}(\boldsymbol{c}))\leq c_{i}(\mathcal{M}_{i}(c_{i}^{\prime},c_{-i}))$).
@@ 3 Indivisible Items with Two Agents
In this section, we assume that there are $n=2$ agents, and items are indivisible. To reduce the task of allocating chores to that of allocating goods, given an arbitrary mechanism $\mathcal{M}^{g}$ (possibly with desirable guarantees for goods), we define a new mechanism $\mathcal{M}^{c}$ such that for every cost profile $\boldsymbol{c}$, $\mathcal{M}^{c}(\boldsymbol{c})=(\mathcal{M}^{g}_{2}(\boldsymbol{c}),\mathcal{M}^{g}_{1}(\boldsymbol{c}))$. In other words, the allocation outputted by $\mathcal{M}^{c}$ results from swapping the bundles in the allocation produced by $\mathcal{M}^{g}$ under the same input. Notice that for all cost profile $\boldsymbol{c}$ and $i\in N$, it holds that $\mathcal{M}^{c}_{i}(\boldsymbol{c})=O\setminus\mathcal{M}^{g}_{i}(\boldsymbol{c})$.
@@
$c_{i}(\mathcal{M}_{i}^{c}(c_{i}^{\prime},c_{3-i}))$ $=c_{i}(O)-c_{i}(\mathcal{M}^{g}_{i}(c_{i}^{\prime},c_{3-i}))$
@@
EF1
Suppose that $\mathcal{M}^{g}$ is EF1 for goods, which implies that for every $i\in N$ with $\mathcal{M}^{g}_{3-i}(\boldsymbol{c})\neq\emptyset$, there exists $o\in\mathcal{M}^{g}_{3-i}(\boldsymbol{c})$ such that $c_{i}(\mathcal{M}^{g}_{i}(\boldsymbol{c}))\geq c_{i}(\mathcal{M}^{g}_{3-i}(\boldsymbol{c})-o)$. Since $o\in\mathcal{M}^{g}_{3-i}(\boldsymbol{c})=\mathcal{M}^{c}_{i}(\boldsymbol{c})$,
$c_{i}(\mathcal{M}_{i}^{c}(\boldsymbol{c})-o)=c_{i}(\mathcal{M}^{g}_{3-i}(\boldsymbol{c})-o)\leq c_{i}(\mathcal{M}_{i}^{g}(\boldsymbol{c}))=c_{i}(\mathcal{M}_{3-i}^{c}(\boldsymbol{c})),$
@@
We give a few remarks here. Firstly, the idea of swapping bundles to connect good and chore allocations with two agents also appears in [BHS20, Seg20] under the contexts of cake cutting and competitive equilibrium, respectively. However, they only leverage this idea in specific settings without further applying it. Moreover, one can similarly establish the connections for properties other than those listed in Theorem 3.1 (e.g., EFX and PO), and we do not pursue this direction for the ease of presentation. Finally, we can analogously reduce the task of allocating goods to that of allocating chores while preserving various properties.
@@ 3.1 Applications
4 Divisible Items
In this section, we consider divisible items and assume that there are $n\geq 2$ agents. We first generalize the reduction given by the previous section to more than two agents. Given an arbitrary mechanism $\mathcal{M}^{g}$ (possibly with desirable guarantees for goods), we define a new mechanism $\mathcal{M}^{c}$ as follows. Given as input a cost profile $\boldsymbol{c}$, let $\mathbf{x}^{g}=\mathcal{M}^{g}(\boldsymbol{c})$, and the output of $\mathcal{M}^{c}(\boldsymbol{c})$, denoted as $\mathbf{x}^{c}$, satisfies $x_{i}^{c}(o)=\frac{1-x_{i}^{g}(o)}{n-1}$ for all $i\in N$ and $o\in O$. To see that $\mathbf{x}^{c}$ defined above is a feasible allocation, notice that for all $i\in N$ and $o\in O$, $x_{i}^{c}(o)\in[0,1]$ since $x_{i}^{g}(o)\in[0,1]$. Moreover, for every $o\in O$,
@@ 4 Divisible Items
Let $\mathcal{M}^{g}$ be a mechanism for divisible items. The following statements hold:
@@
We mention a few direct applications of Theorem 4.1. Recall that [FWVP23] establish an equivalence between the class of fair allocation mechanisms and the class of weakly budget-balanced wagering mechanisms, which results in a family of (strictly) truthful [^4] Informally, a mechanism satisfies strict truthfulness if an agent becomes strictly unhappier when he strictly deviates from his true preference., PROP, and EF mechanisms for goods. Moreover, [AY14] propose the Constrained Serial Dictatorship mechanism, which is truthful and PROP for goods. Notice that the proof of truthfulness in Theorem 4.1 can be easily extended to show that strict truthfulness is also preserved. Therefore, we can apply Theorem 4.1 to the aforementioned mechanisms, leading to new mechanisms with the same (strict) truthfulness and fairness guarantees for chores.

### v2

@@ Abstract
For indivisible chores with two agents, by leveraging the observation that a simple bundle-swapping operation transforms several properties for goods including truthfulness to the corresponding properties for chores, we characterize truthful mechanisms and derive tight guarantees of various fairness notions achieved by truthful mechanisms. Moreover, for homogeneous divisible chores, by generalizing the above transformation to an arbitrary number of agents, we characterize truthful mechanisms with two agents, show that every truthful mechanism with two agents admits an efficiency ratio of 00, and derive a large family of strictly truthful, envy-free (EF), and proportional mechanisms for an arbitrary number of agents. Finally, for indivisible chores with an arbitrary number of agents having bi-valued cost functions, we give an ex-ante truthful, ex-ante Pareto optimal, ex-ante EF, and ex-post envy-free up to one item mechanism, improving the best guarantees for bi-valued instances by prior works.
@@ 1 Introduction
When the goods are indivisible, [ABCM17] characterize the entire family of truthful mechanisms for two agents with additive valuations, which implies that truthfulness is incompatible with any non-trivial fairness notion. Nevertheless, this impossibility result can be bypassed by studying binary valuations [HPPS20, BEF21, BV22], allowing monetary transfer [GIK+24], or resorting to randomized mechanisms [BT24].
For (homogeneous) divisible goods, [CGG13a] give a truthful mechanism that achieves a $1/e$-approximation of Nash welfare with some goods unallocated. More recently, [FWVP23] characterize the set of truthful mechanisms for two agents (subject to some mild technical conditions) and provide a large family of strictly truthful, PROP, and EF mechanisms.
@@
In stark contrast to the fruitful outcomes for truthful good allocation, our understanding of truthful chore allocation still remains considerably more limited. For indivisible chores, [ALW24] study a randomized mechanism that is truthful in expectation and achieve an expected approximation ratio of $O(\sqrt{\log n})$ for MMS. Moreover, [SC25] give a fair and efficient randomized mechanism for restricted additive valuations [^3] Restricted additive valuations constitute a strict superset of binary valuations and a strict subset of additive valuations., whose truthfulness also holds in expectation. To the best of our knowledge, no non-trivial results are known for homogeneous divisible chores.
@@ 1.1 Our Contributions
For indivisible items with two agents, it has been observed by several prior works [BHS20, Seg20, BESS25] that the allocation of goods and that for chores can be equated by the following simple reduction: When allocating goods (resp. chores), we treat the utility (resp. cost) functions as cost (resp. utility) functions and allocate the items as if they are chores (resp. goods), and then we swap the bundles in the allocation. We show that this reduction also preserves truthfulness and various fairness properties (Theorem 3.1). This, together with the characterization by [ABCM17] for goods, allows us to characterize the family of truthful mechanisms for chores with two agents (Theorem 3.2) and derive the tight guarantees achieved by truthful mechanisms with chores for various fairness notions (Corollary 3.3 and Corollary 3.4).
@@
Moving to homogeneous divisible items, we notice that the above reduction from chore allocation to good allocation with two agents can be generalized to an arbitrary number of agents. To illustrate, suppose that we are given a mechanism with prominent properties for goods, and we invoke it with agents’ cost functions as input to obtain an allocation. Intuitively, in this allocation, each agent’s bundle consists of items that are highly undesired by him, and hence, we uniformly distribute each agent’s bundle among all other agents. We show that such an operation transforms fairness and truthfulness for goods to the corresponding properties for chores (Theorem 4.1), which, combined with the results by [FWVP23], enables us to characterize the family of truthful mechanisms for chores with two agents subject to some mild technical conditions (Theorem 4.6) and show that the efficiency ratio, the approximation ratio in social welfare, of every truthful mechanism for chores with two agents is 00 (Theorem 4.7). Moreover, the family of strictly truthful, PROP, and EF mechanisms for goods with an arbitrary number of agents given by [FWVP23] can be easily adapted to mechanisms for chores with the same guarantees (see the remarks after Theorem 4.1).
@@
Finally, we study indivisible items for an arbitrary number of agents with bi-valued cost functions, i.e., the cost of each agent for each item is either $p$ or $q$ for some $p>q>0$, and the bi-valued setting has received extensive attention in recent years [ABF+21a, GM23, EPS22, GMQ22]. Recall that a randomized mechanism satisfies a property ex-ante if it holds in expectation, and satisfies a property ex-post if it holds for every realized integral allocation. We give a randomized mechanism that satisfies ex-ante truthfulness, ex-ante Pareto optimality (PO), ex-ante EF, and ex-post EF1 for chores (Theorem 5.1). This strengthens the guarantees of mechanisms by prior works [EPS22, GMQ22] for the same setting, which only satisfy ex-post fractional PO (fPO) and ex-post EF1 for chores. [^4] Note that ex-ante PO implies ex-post fPO. To achieve this, we leverage the mechanism for homogeneous divisible goods with bi-valued agents recently introduced by [BT24] and the implementation scheme of [Azi20]. In particular, we largely exploit the balance property of the allocations returned by both of the aforementioned components, and considerable efforts are needed to prove the promised guarantees of our mechanism.
@@ 1.2 Other Related Work
During the past decades, there have been plentiful results in the area of fair allocation. In this subsection, we only focus on the most related works in the strategic setting and refer to the recent survey [AAB+23] for a more comprehensive overview of the fair allocation literature.
@@
Besides the cases where resources are modeled as indivisible or homogeneous divisible items, another popular model assumes resources to be a heterogeneous divisible item, which is often referred to as cake cutting. The majority of literature on cake cutting assumes the cake to be a good. The first truthful and EF cake-cutting mechanism for piecewise-uniform valuations is given by [CLPP13], which is later shown to be equivalent to the Maximum Nash Welfare mechanism and satisfy a stronger notion of group strategyproofness [AY14]. For two agents with piecewise-uniform valuations, [BHS20] give a truthful and EF mechanism without the free-disposal assumption and derive the same result when the cake is a chore by the reduction from chores allocation to goods allocation. For the more general piecewise-constant valuations, [BST23] show the non-existence of truthful and PROP cake-cutting mechanisms even for two agents.
@@ 2 Preliminaries
Let $N=[n]$ denote the set of agents and $O=\{o_{1},\ldots,o_{m}\}$ denote the set of items. We will assume that either all items are goods or all of them are chores, and use utilities and costs to respectively describe the values of items. In particular, goods bring non-negative utilities, and agents are happier with a higher utility; conversely, chores bring non-negative costs, and agents are happier with a lower cost. We define the settings of indivisible items and homogeneous divisible items separately.
@@ 2.1 Indivisible Items
2.2 Homogeneous Divisible Items
For homogeneous divisible items, a (fractional) allocation $\mathbf{x}=(x_{1},\ldots,x_{n})\in[0,1]^{n\times m}$ satisfies $\sum_{i\in N}x_{i}(o)=1$ for every $o\in O$, where $x_{i}$ denotes the bundle received by agent $i$ and $x_{i}(o)$ denotes the fraction of item $o$ allocated to agent $i$. When items are goods, each agent $i$ is associated with a non-negative utility function $v_{i}$ where $v_{i}(x_{i})$ represents the utility of agent $i$ for every bundle $x_{i}$. For every $S\subseteq O$, we slightly abuse notations and use $v_{i}(S)$ to denote $v_{i}(\mathbf{1}_{S})$, where $\mathbf{1}_{S}\in[0,1]^{m}$ is the bundle such that $\mathbf{1}_{S}(o)=1$ for every $o\in S$ and $\mathbf{1}_{S}(o)=0$ for every $o\in O\setminus S$. For each $o\in O$, we use $v_{i}(o)$ to denote $v_{i}(\mathbf{1}_{\{o\}})$. We assume utility functions to be additive, i.e., $v_{i}(x_{i})=\sum_{o\in O}x_{i}(o)\cdot v_{i}(o)$. We normalize agents’ utility functions so that $v_{i}(O)=1$ for every $i\in N$. We define the cost functions $c_{i}(\cdot)$ in the same manner for chores. Given an allocation $\mathbf{x}$, the utility (resp. cost) of each agent $i$ is defined as $v_{i}(x_{i})$ (resp. $c_{i}(x_{i})$).
@@ 2.3 Mechanisms
A mechanism $\mathcal{M}$ is truthful for goods (resp. chores) if for every utility profile $\mathbf{v}=(v_{1},\ldots,v_{n})$ (resp. cost profile $\boldsymbol{c}=(c_{1},\ldots,c_{n})$), for all $i\in N$ and utility function $v_{i}^{\prime}$ (resp. cost function $c_{i}^{\prime}$), $v_{i}(\mathcal{M}_{i}(\mathbf{v}))\geq v_{i}(\mathcal{M}_{i}(v_{i}^{\prime},v_{-i}))$ (resp. $c_{i}(\mathcal{M}_{i}(\boldsymbol{c}))\leq c_{i}(\mathcal{M}_{i}(c_{i}^{\prime},c_{-i}))$). We say that $\mathcal{M}$ is strictly truthful if the above inequality is strict whenever $v_{i}^{\prime}\neq v_{i}$.
@@ 3 Indivisible Items with Two Agents
In this section, we assume that there are $n=2$ agents, and items are indivisible. We reduce the task of allocating chores to that of goods via a simple observation in prior works [BHS20, Seg20, BESS25]: Given an arbitrary mechanism $\mathcal{M}^{g}$ (possibly with desirable guarantees for goods), we define a new mechanism $\mathcal{M}^{c}$ such that for every cost profile $\boldsymbol{c}$, $\mathcal{M}^{c}(\boldsymbol{c})=(\mathcal{M}^{g}_{2}(\boldsymbol{c}),\mathcal{M}^{g}_{1}(\boldsymbol{c}))$. In other words, the allocation outputted by $\mathcal{M}^{c}$ results from swapping the bundles in the allocation produced by $\mathcal{M}^{g}$ under the same input. Notice that for all cost profile $\boldsymbol{c}$ and $i\in N$, it holds that $\mathcal{M}^{c}_{i}(\boldsymbol{c})=O\setminus\mathcal{M}^{g}_{i}(\boldsymbol{c})$.
@@
$c_{i}(\mathcal{M}_{i}^{c}(c_{i}^{\prime},c_{-i}))$ $=c_{i}(O)-c_{i}(\mathcal{M}^{g}_{i}(c_{i}^{\prime},c_{-i}))$
@@
EF1.
Suppose that $\mathcal{M}^{g}$ is EF1 for goods, which implies that for every $i\in N$ with $\mathcal{M}^{g}_{-i}(\boldsymbol{c})\neq\emptyset$, there exists $o\in\mathcal{M}^{g}_{-i}(\boldsymbol{c})$ such that $c_{i}(\mathcal{M}^{g}_{i}(\boldsymbol{c}))\geq c_{i}(\mathcal{M}^{g}_{-i}(\boldsymbol{c})-o)$. Since $o\in\mathcal{M}^{g}_{-i}(\boldsymbol{c})=\mathcal{M}^{c}_{i}(\boldsymbol{c})$,
$c_{i}(\mathcal{M}_{i}^{c}(\boldsymbol{c})-o)=c_{i}(\mathcal{M}^{g}_{-i}(\boldsymbol{c})-o)\leq c_{i}(\mathcal{M}_{i}^{g}(\boldsymbol{c}))=c_{i}(\mathcal{M}_{-i}^{c}(\boldsymbol{c})),$
@@
We note here that one can similarly establish the connections for properties other than those listed in Theorem 3.1 (e.g., EFX and PO), and we do not pursue this direction for the ease of presentation. Moreover, we can analogously reduce the task of allocating goods to that of allocating chores while preserving various properties.
@@ 3.1 Applications
4 Homogeneous Divisible Items
In this section, we consider homogeneous divisible items and assume that there are $n\geq 2$ agents. We first generalize the reduction given by the previous section to more than two agents. Given an arbitrary mechanism $\mathcal{M}^{g}$ (possibly with desirable guarantees for goods), we define a new mechanism $\mathcal{M}^{c}$ as follows. Given as input a cost profile $\boldsymbol{c}$, let $\mathbf{x}^{g}=\mathcal{M}^{g}(\boldsymbol{c})$, and the output of $\mathcal{M}^{c}(\boldsymbol{c})$, denoted as $\mathbf{x}^{c}$, satisfies $x_{i}^{c}(o)=\frac{1-x_{i}^{g}(o)}{n-1}$ for all $i\in N$ and $o\in O$. To see that $\mathbf{x}^{c}$ defined above is a feasible allocation, notice that for all $i\in N$ and $o\in O$, $x_{i}^{c}(o)\in[0,1]$ since $x_{i}^{g}(o)\in[0,1]$. Moreover, for every $o\in O$,
@@ 4 Homogeneous Divisible Items
Let $\mathcal{M}^{g}$ be a mechanism for homogeneous divisible items. The following statements hold:
@@
We mention a few direct applications of Theorem 4.1. Recall that [FWVP23] establish an equivalence between the class of fair allocation mechanisms and the class of weakly budget-balanced wagering mechanisms, which results in a family of (strictly) truthful, PROP, and EF mechanisms for goods. Moreover, [AY14] propose the Constrained Serial Dictatorship mechanism, which is truthful and PROP for goods. Notice that the proof of truthfulness in Theorem 4.1 can be easily extended to show that strict truthfulness is also preserved. Therefore, we can apply Theorem 4.1 to the aforementioned mechanisms, leading to new mechanisms with the same (strict) truthfulness and fairness guarantees for chores.